will you please help me with this: what is (12a+6/5a^2)(6/a+1)(5a+5/2a+1)
Answer by fractalier(2101)

In this type of problem, we FACTOR and CANCEL, so we get
(12a+6/5a^2)*(6/a+1)*(5a+5/2a+1) =
6(2a+1)/5a^2)*(6/a+1)*(5(a+1)/2a+1) =
Now the 2a+1 cancels, as do the a+1...that leaves
6/5a^2 * 6 * 5 =
now 5's cancel...leaving
36/a^2